# QAPyramid: Fine-grained Evaluation of Content Selection for Text Summarization

### v1

@@ Abstract
How to properly conduct human evaluations for text summarization is a longstanding challenge. The Pyramid human evaluation protocol, which assesses content selection by breaking the reference summary into sub-units and verifying their presence in the system summary, has been widely adopted. However, it suffers from a lack of systematicity in the definition and granularity of the sub-units. We address these problems by proposing QAPyramid, which decomposes each reference summary into finer-grained question-answer (QA) pairs according to the QA-SRL framework. We collect QA-SRL annotations for reference summaries from CNN/DM and evaluate 10 summarization systems, resulting in 8.9K QA-level annotations. We show that, compared to Pyramid, QAPyramid provides more systematic and fine-grained content selection evaluation while maintaining high inter-annotator agreement without needing expert annotations. Furthermore, we propose metrics that automate the evaluation pipeline and achieve higher correlations with QAPyramid than other widely adopted metrics, allowing future work to accurately and efficiently benchmark summarization systems. [^1] Our data and code can be found in https://github.com/ZhangShiyue/QAPyramid.
@@ 1 Introduction
Figure 1: An illustration of our QAPyramid protocol and its comparison with the ACU protocol Liu et al. (2023b). Compared to ACU, QAPyramid gives credit to finer-grained correctness like “What did something develop? it” (“it” refers to “the vaccine” in the reference summary). The system summary is generated by PEGASUS Zhang et al. (2019). The example is from Liu et al. (2023b) and summaries are truncated due to space limit.
Human evaluation is considered the gold standard for benchmarking progress in text summarization Kryscinski et al. (2019); Bhandari et al. (2020); Fabbri et al. (2021); Celikyilmaz et al. (2021); Krishna et al. (2023), and provides the necessary training or evaluation signals for developing automatic metrics Wei and Jia (2021); Deutsch et al. (2021); Clark et al. (2021). However, there is no consensus on how human evaluation should be conducted. Flawed human evaluation protocol can undermine the reliability of any subsequent automatic evaluations or their outcomes. A key indicator of an unreliable human evaluation is the low inter-annotator agreement, which makes the evaluation results difficult to reproduce.
To make human evaluation more reliable, the Pyramid method Nenkova and Passonneau (2004) was introduced as a reference-based and decomposition-based human evaluation protocol, which is proved to be more reproducible compared to direct assessment (refer to Section 5 for discussions of why reference-based and decomposition-based evaluations are desirable). This method involves decomposing the reference summary into Semantic Content Units (SCUs), defined as “semantically motivated, subsentential units,” and then assessing whether each unit is present (semantically entailed) in the system-generated summary. The protocol has been continuously refined for efficiency and reproducibility Shapira et al. (2019); Bhandari et al. (2020); Zhang and Bansal (2021); Liu et al. (2023b). Despite the widespread recognition, we identify three significant issues with the underlying sub-unit decomposition step, on which the method is based.
@@
To address these problems, we introduce QAPyramid. Our method replaces the SCU generation step with a QA generation step that decomposes the reference summary into Question-Answer pairs (QAs) following the Question-Answer driven Semantic Role Labeling schema (QA-SRL He et al., 2015). Specifically, for each predicate (usually a verb) in the reference summary, annotators generate QA pairs, each corresponding to a “minimal” predicate-argument relation. In Figure 1, five predicates are identified from the reference summary and QAs are generated for each predicate. Then, annotators judge whether each QA is present () or not present () in the system summary and the final score is the number of present QAs over the total number of QAs. Figure 1 also illustrates the comparison between QAPyramid and ACU. QAPyramid demonstrates improvements over Pyramid-style methods for both systematicity and granularity. It not only provides a clear and consistent definition of content units as predicate-argument relations based on QA-SRL, but also decomposes reference summaries into finer-grained units, leading to a more precise score. In addition, QA-SRL annotations are attainable in high quality via crowdsourcing  FitzGerald et al. (2018); Roit et al. (2020); Klein et al. (2020); Brook Weiss et al. (2021), which makes it more scalable compared to expert-annotated SCUs Nenkova and Passonneau (2004); Bhandari et al. (2020) or ACUs Liu et al. (2023b).
We collect QAPyramid annotations on 500 English CNN/DM Hermann et al. (2015) examples. First, we ask crowdsourced workers to write QA pairs for the reference summaries, resulting in 8,652 QA pairs in total. On average, there are 17 QAs (compared to 11 ACUs) per reference summary. Then, across a subset of 50 examples [^2] Due to budget limit, we only manually evaluate systems on a 50-example subset out of the 500. But the complete QAs of 500 examples are still valuable for evaluating systems on a larger set using our automated metrics in the future. and 10 summarization systems, we collect 8,910 manual QA presence judgments. Our analysis reveals that in $21\%$ of the cases, only a subset of the QA pairs for one predicate is present in the system summary, indicating the need for such finer-grained representation to capture partial correctness. Due to QA-SRL formalization, QAPyramid achieves high inter-annotator agreement and a high approval rate for generating QA pairs. And, despite requiring more granular judgments, QAPyramid attains an agreement level in detecting QA presence that is comparable to Pyramid-style human evaluation protocols.
@@
2 Preliminaries
2.1 Pyramid Protocol
Nenkova and Passonneau (2004) proposed the Pyramid protocol to reduce subjectivity and make the human evaluation of content selection more reliable and reproducible. To do so, the protocol is split into two steps. The first step is SCU Generation, in which experts extract fine-grained units, Semantic Content Units (SCUs), from the reference summary. We note that this step only needs to be performed once for each dataset and SCUs can be re-used for evaluating different models’ generations. The second step is Presence Detection (or System Evaluation) which requires annotators to judge whether the system summary contains each SCU as a binary judgment.
Recently, Liu et al. (2023b) revisited the protocol and refined the definition of the sub-unit into atomic content units (ACU): “Elementary information units in the reference summaries, which no longer need to be further split for the purpose of reducing ambiguity in human evaluation.” However, as previously noted, this definition remains ambiguous. In response, we propose a new formalization using QA-SRL.
2.2 QA-SRL
@@
Semantic role labeling (SRL) is to discover the predicate-argument structure of a sentence, i.e., to determine “who does what to whom, when, and where,” etc. Classic SRLs, e.g., FrameNet Baker et al. (1998) and PropBank Palmer et al. (2005), have rather complicated task definitions and require linguistic expertise to conduct annotations. He et al. (2015) introduced QA-SRL to present predicate-argument structure by question-answer (QA) pairs. Without predefining frames or semantic roles, the questions themselves define the set of possible roles. They showed that QA-SRL leads to high-quality SRL annotations and makes scalable data collection possible for annotators with little training and no linguistic expertise, which was later proved by the crowdsourced QA-SRL Bank 2.0 dataset FitzGerald et al. (2018).
The QA-SRL representation has since been utilized in various NLP tasks, demonstrating its versatility and effectiveness. It has been successfully employed for aligning information across different texts Brook Weiss et al. (2021), detecting analogies between procedural texts Sultan and Shahaf (2022), enhancing language modeling pretraining objectives Caciularu et al. (2023) and localizing factual inconsistencies (Cattan et al., 2024). Furthermore, the QA-SRL framework has been successfully extended to capture a wider range of semantic relations, including those found in deverbal nominalizations Klein et al. (2020), discourse relations Pyatkin et al. (2020), and adjectival semantics Pesahov et al. (2023).
@@ 3 QAPyramid: Method and Dataset
Similar to Pyramid, our QAPyramid protocol is composed of two steps: (1) QA generation: write QA pairs for each predicate in the reference summary; (2) QA presence detection: judge the presence of each QA pair in the system summary.
@@ 3.1 QA Generation
For any given dataset, QA generation only needs to be done once for the reference summaries in its evaluation set. We choose CNN/DM Hermann et al. (2015), one of the most popularly used text summarization datasets, as our test bed. We use a subset of 500 examples of the CNN/DM test set, the same set used by Liu et al. (2023b). For each reference summary, we first extract predicates automatically via AllenNLP SRL API Gardner et al. (2018) which uses spaCy under the hood. On average, each reference contains 7.6 predicates.
Then, for each predicate, we ask one human annotator to write QA pairs. Figure 6 in the Appendix is our annotation UI on Amazon Mechanical Turk for writing QA pairs. Note that the previous QA-SRL annotation FitzGerald et al. (2018) requires annotators to follow predefined automata, which sometimes results in non-natural questions. Here we opt for a less confined approach to allow annotators to write questions following our instructions (instructions can be seen in Figure 6). We show one sentence at a time and highlight one predicate in the sentence. The annotator is asked to write at most 5 questions and, for each question, fill in at most 3 answers. To recruit high-quality annotators for this task, we initially picked 4 examples as qualification tasks. Workers were qualified only if they correctly wrote QA pairs for all 4 tasks. Eventually, we recruited 31 annotators.
@@
After completing the data collection, we identified two issues. First, approximately 10% of the data contained duplicated QA pairs—that is, questions and answers with the same meaning but different wording—for the same predicate within a single reference. We (the authors) manually de-duplicated these instances; this deduplication task can be easily integrated into the QA verification step in future work. Second, about 10 summaries had fewer than five QA pairs. This was mainly because no predicates were extracted from sentences lacking verbs, such as “Lewis Hamilton over half-a-second clear of team-mate Nico Rosberg.” To address this, we manually write the QA pairs for these examples. Ultimately, we compiled a total of 8,652 QA pairs, averaging 17 QAs per reference summary.On the same dataset, there are on average 11 ACUs Liu et al. (2023b) per reference summary, which confirms the finer granularity of our evaluation protocol.
@@ 3.2 QA Presence Detection
After we obtained all the QA pairs, the next step is to conduct system evaluations. For any system summary, we ask humans to judge whether each QA is present () or not present () in the system summary. Figure 8 in the Appendix shows the annotation UIs of this task. We instruct annotators to judge a QA pair as being present if its meaning is covered or implied by the system summary, or can be inferred from the system summary, while the exact format, expression, or wording can be different. Note that the reference summary is also provided in this task for annotators to ground each QA and infer any necessary coreference information, e.g., in Figure 1, we can easily know “it” refers to “the vaccine” based on the reference. Same as QA generation and verification tasks, we used 4 HITs as qualification tasks, and only employed annotators who got more than 90% accuracy. We recruited 27 annotators in the end. To reduce workload, we only display the QA pairs for one predicate (usually 2 or 3 QAs) in one HIT and pay \$0.2/HIT, which makes a \$10 hourly rate.
Due to budget constraints, we subsampled 50 examples from the 500 CNN/DM test examples to conduct system evaluation and evaluated 10 systems, including 5 models that are fine-tuned on CNN/DM: BART Lewis et al. (2020), PEGASUS Zhang et al. (2019), BRIO Liu et al. (2022), BRIO-Ext, and MatchSum Zhong et al. (2020), and 5 LLMs that generate summaries via 1-shot learning: Llama-3-8B-Instruct Dubey et al. (2024), Llama-3-70B-Instruct, Mixtral-8$\times$7B-Instruct Jiang et al. (2024), Mixtral-8$\times$22B-Instruct, and GPT-4 (gpt-4-0125). In total, we collected 8,910 QA presence judgments. Each judgment was provided independently by three annotators, and the final decision was determined by a majority vote. The average inter-annotator agreement (IAA) is 0.74 (Krippendorff’s alpha). For reference, Liu et al. (2023b) reported an IAA of 0.75, Zhang and Bansal (2021) reported 0.73, and Bhandari et al. (2020) reported 0.66.
QAPyramid $n$QAPyramid ACU $n$ACU R2-R R2-F1 Length
@@
Mixtral-8$\times$7B-Instruct 0.48 0.45 - - 0.19 0.17 67.44
Mixtral-8$\times$22B-Instruct 0.48 0.44 - - 0.23 0.18 74.58
@@
Table 1: The evaluation results of different summarization systems on 50 CNN/DM testing examples. QAPyramid and $n$QAPyramid are unnormalized and normalized QAPyramid scores. ACU and $n$ACU are unnormalized and normalized ACU scores Liu et al. (2023b). R2-R and R2-F1 are ROUGE-2 recall and f1 scores Lin (2004). Length is the average summary length in tokens. Table 6 contains systems scores of many other metrics. FT means the models fine-tuned on CNN/DM training set, and LLM means the 1-shot LLM-based models
@@ 3.3 Summary Scoring
Given human annotations, we now can calculate the QAPyramid scores. For any given system summary $s_{i}$, assume its corresponding reference summary $r_{i}$ has $K_{i}$ QA pairs $\{\text{QA}_{ij}\}_{j=1}^{K_{i}}$, then QAPyramid is defined as the number of present QA pairs in $s_{i}$ divided by $K_{i}$:
@@
This metric is essentially a recall, i.e., a longer summary with more information usually gets a higher score. In an extreme case, when the summary is the same as the source document, it receives a perfect score. To combat this issue, Liu et al. (2023b) introduced a normalized ACU score, $n$ACU, by multiplying the original ACU score by a length penalty, i.e., $n$ACU $=p^{l}_{i}*$ACU, where $p^{l}_{i}=e^{\min(0,\frac{1-\frac{|s_{i}|}{|r_{i}|}}{\alpha})}$ and $|s_{i}|,|r_{i}|$ are the length (i.e., number of words) of the system and reference summary. Essentially, system summaries that are longer than the reference summary get discounted scores. In practice, $\alpha$ is set by de-correlating $n$ACU with summary length.
However, this length penalty, especially how $\alpha$ is being set, assumes the unnormalized score is positively correlated with summary length, i.e., longer summaries tend to get higher scores. This is usually true for fine-tuned models. But for non-finetuned LLMs, they sometimes suffer from the degeneration problem Holtzman et al. (2020) – getting stuck into a repetition loop and generating a sentence or a sub-sentence repeatedly, see a degenerated summary produced by Llama-3-8B-Instruct in Figure 2. In this case, the extending length does not provide more information and thus does not lead to a higher score. Though this extending length also needs to be penalized, it has unique behavior and needs to be dealt with separately. Therefore, we introduce a novel repetition penalty $p_{i}^{r}=1-rp_{i}$, where $rp_{i}$ is the repetition rate of $s_{i}$ meaning what percentage of the summary is repetition (please refer to Figure 2 in Appendix A.1). Then we change the length penalty to $p^{l}_{i}=e^{\min(0,\frac{1-\frac{|s_{i}|*p_{i}^{r}}{|r_{i}|}}{\alpha})}$, where $|s_{i}|*p_{i}^{r}$ is the effective summary length – the length of non-repetitive text. We set $\alpha$ by de-correlating $p^{l}_{i}*\text{QAPyramid{}}$ with effective summary length. Using our collected data, the Pearson correlation between effective summary length and QAPyramid is 0.27. After setting $\alpha=6$, the correlation between effective summary length and $p^{l}_{i}*\text{QAPyramid{}}$ reduced to -0.01 (see details in Table 5 in Appendix A.1).

### v2

@@ Abstract
How to properly conduct human evaluations for text summarization is a longstanding challenge. The Pyramid human evaluation protocol, which assesses content selection by breaking the reference summary into sub-units and verifying their presence in the system summary, has been widely adopted. However, it suffers from a lack of systematicity in the definition and granularity of the sub-units. We address these problems by proposing QAPyramid, which decomposes each reference summary into finer-grained question-answer (QA) pairs according to the QA-SRL framework. We collect QA-SRL annotations for reference summaries from CNN/DM and evaluate 10 summarization systems, resulting in 8.9K QA-level annotations. We show that, compared to Pyramid, QAPyramid provides more systematic and fine-grained content selection evaluation while maintaining high inter-annotator agreement without needing expert annotations. Furthermore, we propose metrics that automate the evaluation pipeline and achieve higher correlations with QAPyramid than other widely adopted metrics. [^1] Our data and code can be found in https://github.com/ZhangShiyue/QAPyramid.
@@ 1 Introduction
Figure 1: An illustration of our QAPyramid protocol and its comparison with the ACU protocol (Liu et al., 2023b). Compared to ACU, QAPyramid gives credit to finer-grained correctness like “What did something develop? it” (“it” refers to “the vaccine” in the reference summary). The system summary is generated by PEGASUS (Zhang et al., 2019). The example is from Liu et al. (2023b) and summaries are truncated due to space limit.
Human evaluation is considered the gold standard for benchmarking progress in text summarization (Kryscinski et al., 2019; Bhandari et al., 2020; Fabbri et al., 2021b; Celikyilmaz et al., 2021; Krishna et al., 2023), and provides the necessary training or evaluation signals for developing automatic metrics (Wei & Jia, 2021; Deutsch et al., 2021; Clark et al., 2021). However, there is no consensus on how human evaluation should be conducted. Flawed human evaluation protocol can undermine the reliability of any subsequent automatic evaluations or their outcomes. A key indicator of an unreliable human evaluation is the low inter-annotator agreement, which makes the evaluation results difficult to reproduce.
To make human evaluation more reliable, the Pyramid method (Nenkova & Passonneau, 2004) was introduced as a reference-based and decomposition-based human evaluation protocol. The reference defines what content should be selected for the summary, and decomposition reduces ambiguity when evaluating a long summary. Because of these, Pyramid is proved to be more reproducible compared to direct assessment (see more discussions in Section 2). In practice, Pyramid first decomposes the reference summary into Semantic Content Units (SCUs), defined as “semantically motivated, subsentential units,” and then assesses whether each unit is present (semantically entailed) in the system-generated summary. The protocol has been continuously refined for efficiency and reproducibility (Shapira et al., 2019; Bhandari et al., 2020; Zhang & Bansal, 2021; Liu et al., 2023b). Despite the widespread recognition, we identify three significant issues with the underlying sub-unit decomposition step, on which the method is based.
@@
To address these problems, we introduce QAPyramid. Our method replaces the SCU generation step with a QA generation step that decomposes the reference summary into Question-Answer pairs (QAs) following the Question-Answer driven Semantic Role Labeling schema (QA-SRL He et al., 2015). Specifically, for each predicate (usually a verb) in the reference summary, annotators generate QA pairs, each corresponding to a “minimal” predicate-argument relation. In Figure 1, five predicates are identified from the reference summary and QAs are generated for each predicate. Then, annotators judge whether each QA is present () or not present () in the system summary and the final score is the number of present QAs over the total number of QAs. Figure 1 also illustrates the comparison between QAPyramid and ACU. QAPyramid demonstrates improvements over Pyramid-style methods for both systematicity and granularity. It not only provides a clear and consistent definition of content units as predicate-argument relations based on QA-SRL, but also decomposes reference summaries into finer-grained units, leading to a more precise score. In addition, QA-SRL annotations are attainable in high quality via crowdsourcing  (FitzGerald et al., 2018; Roit et al., 2020; Klein et al., 2020; Brook Weiss et al., 2021), which makes it more scalable compared to expert-annotated SCUs (Nenkova & Passonneau, 2004; Bhandari et al., 2020) or ACUs (Liu et al., 2023b).
We collect QAPyramid annotations on 500 English CNN/DM (Hermann et al., 2015) examples. First, we ask crowdsourced workers to write QA pairs for the reference summaries, [^2] Some may challenge the quality of reference summaries in the CNN/DM dataset. However, this problem is orthogonal to our contribution. For any type of reference (good or bad), QAPyramid is more systematic, finer-grained, and more scalable than Pyramid-style protocols. Improving the quality of references is an orthogonal problem to be resolved. resulting in 8,652 QA pairs in total. On average, there are 17 QAs (compared to 11 ACUs) per reference summary. Then, across a subset of 50 examples [^3] Due to budget limit, we only manually evaluate systems on a 50-example subset out of the 500. But the complete QAs of 500 examples are still valuable for evaluating systems on a larger set using our automated metrics in the future. and 10 summarization systems, we collect 8,910 manual QA presence judgments. Our analysis reveals that in $21\%$ of the cases, only a subset of the QA pairs for one predicate is present in the system summary, indicating the need for such finer-grained representation to capture partial correctness. Due to QA-SRL formalization, QAPyramid achieves high inter-annotator agreement and a high approval rate for generating QA pairs. And, despite requiring more granular judgments, QAPyramid attains an agreement level in detecting QA presence that is comparable to Pyramid-style human evaluation protocols.
@@
2 Background and Related Works
Human Evaluation for Text Summarization.
In many cases, human evaluation is conducted via direct rating, i.e., humans directly rate the quality of the summary (or rate for certain aspects, e.g., relevance (Fabbri et al., 2021b)). However, direct rating often suffers from subjectivity, low agreement, and thus non-reproducibility (Kiritchenko & Mohammad, 2017; Falke et al., 2019; Shapira et al., 2019). Two factors contribute to this issue: (1) what content is considered important and should be selected into the summary may vary from person to person, and (2) the summary is usually more than one sentence, and different annotations may put their focus on different parts of the summary. The canonical Pyramid (Nenkova & Passonneau, 2004; Shapira et al., 2019) method addressed these issues using reference summaries and decompositions. Recently, Liu et al. (2023b) revisited Pyramid and refined the definition of the sub-unit into atomic content units (ACU): “Elementary information units in the reference summaries, which no longer need to be further split for the purpose of reducing ambiguity in human evaluation.” However, there are still no standard guidelines on how to write each unit and what content should be included in each unit. Hence, you may get different sets of units from different annotators, which undermines reproducibility. In response, we propose a new formalization using QA-SRL.
QA-SRL.
@@ 2 Background and Related Works
Semantic role labeling (SRL) is to discover the predicate-argument structure of a sentence, i.e., to determine “who does what to whom, when, and where,” etc. Classic SRLs, e.g., FrameNet (Baker et al., 1998) and PropBank (Palmer et al., 2005), have rather complicated task definitions and require linguistic expertise to conduct annotations. He et al. (2015) introduced QA-SRL to present predicate-argument structure by question-answer (QA) pairs. Without predefining frames or semantic roles, the questions themselves define the set of possible roles. They showed that QA-SRL leads to high-quality SRL annotations and makes scalable data collection possible for annotators with little training and no linguistic expertise, which was later proved by the crowdsourced QA-SRL Bank 2.0 dataset (FitzGerald et al., 2018). The QA-SRL representation has since been utilized in various NLP tasks, demonstrating its versatility and effectiveness (Brook Weiss et al., 2021; Sultan & Shahaf, 2022; Caciularu et al., 2023; Cattan et al., 2024).
Automatic Evaluation for Text Summarization.
Compared to human evaluations, automatic metrics are fast, cheap, and reproducible. However, how well they correlate with human judgment is always a concern. Over the years, many automatic evaluation metrics have been introduced. Some early metrics measure the n-gram overlap between system and reference summaries (Papineni et al., 2002; Lin, 2004; Banerjee & Lavie, 2005). To alleviate the rigidity of exact lexical match, metrics based on the similarity between embeddings were introduced (Zhao et al., 2019; Clark et al., 2019; Zhang* et al., 2020). There are also works automating the Pyramid protocol (Yang et al., 2016; Hirao et al., 2018; Gao et al., 2019; Zhang & Bansal, 2021; Liu et al., 2023c; Nawrath et al., 2024). They use Open IE, SRL, AMR, fine-tuned models, or LLMs to automate unit extraction and use NLI models to automate unit presence detection. Compared to these works, our AutoQAPyramid is advantageous because it automates a more systermatical, reproducible, and fine-grained human evaluation protocol, QAPyramid. Besides evaluating content selection, a separate line of research has been focused on faithfulness or factuality evaluation, i.e., checking if the source document(s) entail the summary (Cao et al., 2018; Maynez et al., 2020; Laban et al., 2022; Zha et al., 2023). The Pyramid type of methods have also been applied in this scenario (Chen et al., 2023; Min et al., 2023; Kamoi et al., 2023; Wan et al., 2024; Tang et al., 2024; Gunjal & Durrett, 2024). Some methods are based on question generation and answering (Wang et al., 2020; Durmus et al., 2020; Scialom et al., 2021; Fabbri et al., 2022). However, their QAs are not QA-SRL-based questions, and each QA contains more than one argument.
@@ 3.1 QA Generation
For any given dataset, QA generation only needs to be done once for the reference summaries in its evaluation set. We choose CNN/DM (Hermann et al., 2015), one of the most popularly used text summarization datasets, as our test bed. We use a subset of 500 examples of the CNN/DM test set, the same set used by Liu et al. (2023b). For each reference summary, we first extract predicates automatically via AllenNLP SRL API (Gardner et al., 2018) which uses spaCy under the hood. On average, each reference contains 7.6 predicates.
Then, for each predicate, we ask one human annotator to write QA pairs. Figure 6 in the Appendix is our annotation UI on Amazon Mechanical Turk for writing QA pairs. Note that the previous QA-SRL annotation (FitzGerald et al., 2018) requires annotators to follow predefined automata, which sometimes results in non-natural questions. Here we opt for a less confined approach to allow annotators to write questions following our instructions (instructions can be seen in Figure 6). We show one sentence at a time and highlight one predicate in the sentence. The annotator is asked to write at most 5 questions and, for each question, fill in at most 3 answers. To recruit high-quality annotators for this task, we initially picked 4 examples as qualification tasks. Workers were qualified only if they correctly wrote QA pairs for all 4 tasks. Eventually, we recruited 31 annotators.
@@
Eventually, we collected a total of 8,652 QA pairs, averaging 17 QAs per reference summary. On the same dataset, there are on average 11 ACUs (Liu et al., 2023b) per reference summary, which confirms the finer granularity of our evaluation protocol.
@@ 3.2 QA Presence Detection
After we obtained all the QA pairs, the next step is to conduct system evaluations. For any system summary, we ask humans to judge whether each QA is present () or not present () in the system summary. Figure 8 in the Appendix shows the annotation UIs of this task. We instruct annotators to judge a QA pair as being present if its meaning is covered or implied by the system summary, or can be inferred from the system summary, while the exact format, expression, or wording can be different. Note that the reference summary is also provided in this task for annotators to ground each QA and infer any necessary coreference information, e.g., in Figure 1, we can easily know “it” refers to “the vaccine” based on the reference. Same as QA generation and verification tasks, we used 4 HITs as qualification tasks, and only employed annotators who got more than 90% accuracy. We recruited 27 annotators. To reduce workload, we only display the QA pairs for one predicate (usually 2 or 3 QAs) in one HIT and pay \$0.2/HIT, which makes a \$10 hourly rate.
[…]
QAPyramid nnQAPyramid ACU nnACU R2-R R2-F1 Length
@@
[…]
@@
Table 1: The evaluation results of different summarization systems on 50 CNN/DM testing examples. QAPyramid and nnQAPyramid are unnormalized and normalized QAPyramid scores. ACU and nnACU are unnormalized and normalized ACU scores (Liu et al., 2023b). R2-R and R2-F1 are ROUGE-2 recall and f1 scores (Lin, 2004). Length is the average summary length in tokens. Table 6 contains systems scores of many other metrics. FT means the models fine-tuned on CNN/DM training set, and LLM means the 1-shot LLM-based models
@@ 3.3 Summary Scoring
Given human annotations, we now can calculate the QAPyramid scores. For any given system summary $s_{i}$, assume its reference summary $r_{i}$ has $K_{i}$ QA pairs $\{\text{QA}_{ij}\}_{j=1}^{K_{i}}$, then QAPyramid is defined as the number of present QA pairs in $s_{i}$ divided by $K_{i}$:
@@
This metric is essentially a recall, i.e., a longer summary with more information usually gets a higher score. In an extreme case, when the summary is the same as the source document, it receives a perfect score. To combat this issue, Liu et al. (2023b) introduced a normalized ACU score, nnACU, by multiplying the original ACU score by a length penalty, i.e., nnACU $p^{l}_{i}*$ACU, where $p^{l}_{i}=e^{\min(0,\frac{1-\frac{|s_{i}|}{|r_{i}|}}{\alpha})}$ and $|s_{i}|,|r_{i}|$ are the length (i.e., number of words) of the system and reference summary. Essentially, system summaries that are longer than the reference summary get discounted scores. In practice, $\alpha$ is set by de-correlating nnACU with summary length.
However, this length penalty, especially how $\alpha$ is being set, assumes the unnormalized score is positively correlated with summary length, i.e., longer summaries tend to get higher scores. This is usually true for fine-tuned models. But for non-finetuned LLMs, they sometimes suffer from the degeneration problem (Holtzman et al., 2020) – getting stuck into a repetition loop and generating a sentence or a sub-sentence repeatedly, see a degenerated summary produced by Llama-3-8B-Instruct in Figure 2. In this case, the extending length does not provide more information and thus does not lead to a higher score. Though this extending length also needs to be penalized, it has unique behavior and needs to be dealt with separately. Therefore, we introduce a novel repetition penalty $p_{i}^{r}=1-rp_{i}$, where $rp_{i}$ is the repetition rate of $s_{i}$ meaning what percentage of the summary is repetition (please refer to Figure 2 in Appendix A.2). Then we change the length penalty to $p^{l}_{i}=e^{\min(0,\frac{1-\frac{|s_{i}|*p_{i}^{r}}{|r_{i}|}}{\alpha})}$, where $|s_{i}|*p_{i}^{r}$ is the effective summary length – the length of non-repetitive text. We set $\alpha$ by de-correlating $p^{l}_{i}*\text{QAPyramid{}}$ with effective summary length. Using our collected data, the Pearson correlation between effective summary length and QAPyramid is 0.27. After setting $\alpha=6$, the correlation between effective summary length and $p^{l}_{i}*\text{QAPyramid{}}$ reduced to -0.01 (see details in Table 5 in Appendix A.2).